# Strong Baseline: Multi-UAV Tracking via YOLOv12 with BoT-SORT-ReID

### v1

@@ Abstract
Detecting and tracking multiple unmanned aerial vehicles (UAVs) in thermal infrared video is inherently challenging due to low contrast, environmental noise, and small target sizes. This paper provides a straightforward approach to address multi-UAV tracking in thermal infrared video, leveraging recent advances in detection and tracking. Instead of relying on the YOLOv5 with the DeepSORT pipeline, we present a tracking framework built on YOLOv12 and BoT-SORT, enhanced with tailored training and inference strategies. We evaluate our approach following the metrics from the 4th Anti-UAV Challenge and demonstrate competitive performance. Notably, we achieve strong results without using contrast enhancement or temporal information fusion to enrich UAV features, highlighting our approach as a ”Strong Baseline” for the multi-UAV tracking task. We provide implementation details, in-depth experimental analysis, and a discussion of potential improvements. The code is available at https://github.com/wish44165/YOLOv12-BoT-SORT-ReID.
@@ 1 Introduction
Thermal infrared videos offer advantages over traditional RGB imagery, such as enhanced visibility in low-light and adverse weather conditions, making them ideal for security and surveillance applications. This paper focuses on using thermal infrared video for multi-UAV tracking, exploiting its importance in challenging environments where RGB-based methods may fail. Fig. 1 (a) illustrates thermal infrared frames with diverse backgrounds from the MOT training set, while Fig. 1 (b) highlights minor defects, such as annotation errors, redundancies, missed labels, and low-quality frames, that account for a negligible portion of the dataset and can be safely disregarded during training. Additionally, Fig. 2 displays cropped image patches from bounding box annotations in the training set, illustrating the varying sizes of UAVs, from several pixels to single-digit pixels. We build a complete UAV tracking workflow by leveraging the latest YOLOv12 [36] detector and BoT-SORT [1] tracking algorithm, which outperform the well-established YOLOv5 [18] with DeepSORT [40] combination. We also implement some strategies to enhance multi-UAV tracking performance further. Our contributions are as follows:
@@ 3.1 Problem Statement
The goal is to track UAVs as accurately as possible, with evaluation metrics detailed in Sec. 4.1. The challenge consists of three tracks, each corresponding to a different scenario. Track 1 and Track 2 are SOT tasks, differing in whether the UAV’s initial location is given. Track 3 is a MOT task where the initial locations of UAVs are provided.
@@ 3.2 Data Analysis and Preparation
We first analyze each track’s training and testing data, as summarized in Tab. 1. Track 1 and Track 2 share the same training set, consisting of 23 sequences at 512$\times$512 resolution and 200 sequences at 640$\times$512 resolution. The training set for Track 3 is composed of 200 sequences at 640$\times$512 resolution. For testing, Track 1 and Track 2 each contain 216 non-overlapping sequences. Track 1’s test set is entirely at 640$\times$512 resolution, while Track 2 includes 16 sequences at 640$\times$512 and 200 at 512$\times$512. Track 3’s test set consists of 100 sequences at 640$\times$512 resolution. Additionally, Tab. 1 reports the width, height, and area distributions, along with their mean and standard deviation, providing essential insights for model hyperparameter tuning. Note that there may be slight differences in the numbers compared to the official release, as we have removed redundant annotations and defect cases, as illustrated in Fig. 1 (b).
@@
Table 1: Summary of data characteristics for training and testing. The training data for Track 1 and Track 2 are identical, while the testing data for Track 1 and Track 2 have no overlapping parts. Since Track 2 does not provide initial bounding boxes, we use ”N/A” to indicate that this information is not applicable. Note that some bounding boxes were removed in Track 1 and Track 2 training sets due to labeled non-existent, zero-sized width or height, and cases where the boxes covered the entire image.
After analyzing the data, we split it for model training preparation. The number of frames and bounding boxes used for training, validation, and testing in the SOT and MOT tasks are detailed in Tab. 2. Specifically, Track 1 and Track 2 use YOLOv12 with BoT-SORT, while Track 3 employs YOLOv12 with BoT-SORT-ReID. Note that some numbers are in parentheses since we found the test set to provide limited information for the SOT task. Thus, the values in parentheses reflect the data only split into training and validation sets. Additionally, for BoT-SORT training, 1/10 of the data is primarily used to train the ReID module. This approach provides more effective ReID module training since many scenes are visually similar.
@@
Table 2: Data preparation summary for YOLOv12 and BoT-SORT-ReID across single-object (Track 1 and Track 2) and multi-object (Track 3) tracking. For YOLOv12 in SOT, numbers in parentheses indicate an alternative split with only training and validation sets. For BoT-SORT-ReID, numbers in parentheses indicate a reduced ReID training set selecting only the first 1/10 of frames from each sequence.
@@ 3.4 Training and Inference Strategies
The inference workflow is presented in Fig. 3. The overall procedure follows the original BoT-SORT scheme. However, we modify the output by reporting both online and lost targets for Track 1 and Track 2 while preserving the original output format for Track 3. We did not use linear track interpolation because ID switching frequently occurs due to camera motion or fast-moving UAVs, making interpolation ineffective for recovering missing detections. Instead, for the SOT task, we adopt a strategy based on the assumption that each frame contains at most one UAV, following this priority order: (1) report the UAV with the highest confidence score among the online targets, (2) if no online target is available, continue reporting the previous ID as the lost target in the subsequent tracker buffer frames, (3) if no previous ID is available, report the last known location until new online targets are detected. This strategy leverages the Kalman Filter’s prediction to accurately estimate the UAV’s location based on prior positions and velocity, significantly improving evaluation metrics in the SOT task. However, this strategy is not feasible for the MOT task due to the frequent overlap and ID switching between online and lost targets, which would lead to poor results. Therefore, we maintain the original output for Track 3 in this case.
Figure 3: YOLOv12n with BoT-SORT-SBS-S50 workflow diagram. The workflow follows the original BoT-SORT framework [1], with a slight revision: incorporating lost tracks to compensate for uninformative frames and improve object continuity. Specifically, for Track 1 and Track 2, lost target information is used to annotate potential object locations, while Track 3 retains the BoT-SORT original output.
@@ 4 Experimental Results
The experiments were conducted on two platforms: the first, a system with an Intel Core i7-12650H CPU, an NVIDIA RTX 4050 GPU, and 16 GB of RAM; the second, a high-performance computing (HPC) system [27], equipped with an NVIDIA H100 GPU and 80 GB of memory. All models were trained using the default settings (e.g., image input size of 640 and a track buffer of 30 frames) unless otherwise specified in the content or tables. This section begins by outlining the evaluation metrics for the three tracks, followed by the results for both the SOT and MOT tasks. We then present the leaderboard rankings and discuss key considerations and potential image enhancement techniques that could further improve UAV tracking.
@@ 4.1 Evaluation Metrics
Two evaluation metrics are used across the three competition tracks. The first metric applies to Track 1 and Track 2, where tracking accuracy is defined as:
@@
4.2 Evaluation Results on Track 1 and Track 2
@@
Table 3: Evaluation results for Track 1 and Track 2, summarizing eight trials per track. The first two rows detail YOLOv12 configurations, varying model sizes (n, s, m, l, x), and training epochs. The third row specifies the use of BoT-SORT (”✗” indicates exclusion; ”✓” indicates inclusion). The fourth row lists the ”min box area” threshold for filtering tiny bounding boxes, adjusted to 4 and 1 to account for tiny UAVs. The final row reports the resulting scores.
We present the evaluation results for Track 1 and Track 2 together, as both are SOT tasks, with the only difference being the presence of the initial UAV location. Eight meaningful trials are selected for both tracks, as shown in Tab. 3. Trials 1 and 2 serve as an ablation study to assess the impact of BoT-SORT. The results demonstrate a significant performance improvement: the score in Track 1 increases from 0.0786 to 0.5529, and in Track 2, it rises from 0.0992 to 0.3106 simply by adding BoT-SORT after the YOLOv12n detector. Trials 2 through 6 evaluate different detector model sizes (n, s, m, l, x), with the highest score achieved using YOLOv12l for both tracks. Trial 7 examines the effect of extended 300 epochs training, revealing a decline in performance compared to the 100 epochs training, likely due to overfitting. Finally, Trial 8 for each track shows the highest score we submitted, tuning the minimum box area threshold from 10 to 4 for Track 1 and from 10 to 1 for Track 2 to better capture smaller UAVs that may have been missed with the default setting.
@@ 4.5 Discussion and Enhancement Techniques
Figure 5: Potential frame enhancement techniques for multi-UAV tracking on a MultiUAV-262 video frame. From left to right: (1) original thermal infrared frame, (2) Sobel edge-based sharpening [10], (3) contrast enhancement via Contrast Limited Adaptive Histogram Equalization (CLAHE) [28], and (4) ReynoldsFlow+ visualization highlighting motion patterns to assist UAV detection [4].
@@
The evaluation results reveal several key insights. First, overfitting emerged due to our data-splitting strategy. To maximize scene diversity, we did not categorize videos by attributes such as fixed-camera setups or background types (e.g., sky or buildings). Instead, we directly split the dataset into training, validation, and testing sets, occasionally allowing frames from the same video to appear across splits. This likely contributed to overfitting, as evidenced by AP score discrepancies during local testing. Second, accurately rescaling the provided initial object positions to match the resolution used in training and inference is critical, as mismatches can mislead the tracker and degrade subsequent predictions. Third, increasing image resolution is key to breaking performance plateaus when parameter tuning fails to improve accuracy. For example, scaling from 640 to 1280 resolution yielded a significant score improvement of approximately 0.1. However, further increases produced diminishing gains as training at 2560 pixels for 7 epochs reached a score of 0.7072, and training at 3840 pixels for 1 epoch reached 0.7098, while both required significantly higher computational costs compared to training at 1280 pixels. Fourth, memory consumption during inference with YOLOv12 and BoT-SORT-ReID accumulates over time, leading to program crashes. To address this, we executed inference on a per-folder basis rather than processing all sequences in a single run. Finally, there is a clear performance gap between runs with accurate initial object positions and those without, as reflected in the performance difference between Track 1 and Track 2. This underscores the importance of promptly and reliably estimating initial positions to boost performance further.
@@
Additionally, as previously discussed, while our approach provides a strong baseline, it remains insufficient for achieving top-tier performance without further refinement. Fig. 4 (a) displays our model’s predictions across various scenarios, while Fig. 4 (b) highlights several key failure cases: (1) overlapping UAVs frequently cause ID switches, (2) distinguishing UAVs from flying creatures remains difficult, with the model often reassigning new IDs to UAVs following brief occlusions, (3) complex backgrounds lead to missed detections and tracking failures, and (4) tiny UAVs in cluttered environments provide little to no valuable information, making detection highly unreliable. The corresponding heatmaps in the last row illustrate the model’s inability to perceive UAVs effectively in these challenging conditions. These limitations emphasize the importance of image enhancement techniques to improve performance further. Fig. 5 illustrates several potential image enhancement methods. From left to right: (1) the original thermal infrared frame, (2) Sobel edge-based sharpening [10], which highlights edges more clearly than the original, (3) Contrast Limited Adaptive Histogram Equalization (CLAHE) [28], which improves contrast, and (4) ReynoldsFlow+ [4], a temporal enhancement method based on the Reynolds Transport Theorem [31], a three-dimensional generalization of the Leibniz integral rule [9], providing enhanced appearance for moving UAVs.

### v2

@@ Abstract
Detecting and tracking multiple unmanned aerial vehicles (UAVs) in thermal infrared video is inherently challenging due to low contrast, environmental noise, and small target sizes. This paper provides a straightforward approach to address multi-UAV tracking in thermal infrared video, leveraging recent advances in detection and tracking. Instead of relying on the well-established YOLOv5 with DeepSORT combination, we present a tracking framework built on YOLOv12 and BoT-SORT, enhanced with tailored training and inference strategies. We evaluate our approach following the 4th Anti-UAV Challenge metrics and reach competitive performance. Notably, we achieved strong results without using contrast enhancement or temporal information fusion to enrich UAV features, highlighting our approach as a "Strong Baseline" for multi-UAV tracking tasks. We provide implementation details, in-depth experimental analysis, and a discussion of potential improvements. The code is available at https://github.com/wish44165/YOLOv12-BoT-SORT-ReID.
@@ 1 Introduction
Thermal infrared videos offer advantages over traditional RGB imagery, such as enhanced visibility in low-light and adverse weather conditions, making them ideal for security and surveillance applications. This paper focuses on using thermal infrared video for multi-UAV tracking, exploiting its importance in challenging environments where RGB-based methods may fail. Fig. 1 (a) illustrates thermal infrared frames with diverse backgrounds from the MOT training set, while Fig. 1 (b) highlights minor defects, such as annotation errors, redundancies, missed labels, and low-quality frames, that account for a negligible portion of the dataset and can be safely disregarded during training. Additionally, Fig. 2 displays cropped image patches from bounding box annotations in the training set, illustrating the varying sizes of UAVs, from several pixels to single-digit pixels. We build a complete UAV tracking workflow by leveraging the latest YOLOv12 [36] detector and BoT-SORT [1] tracking algorithm, which outperform the YOLOv5 [18] with the DeepSORT [40] pipeline. We also implement some strategies to enhance multi-UAV tracking performance further. Our contributions are as follows:
@@ 3.1 Problem Statement
The goal is to track UAVs as accurately as possible, with evaluation metrics detailed in Sec. 4.1. The challenge consists of three tracks, each corresponding to a different scenario. Tracks 1 and 2 are SOT tasks, differing in whether the UAV’s initial location is given. Track 3 is a MOT task where the initial locations of UAVs are provided.
@@ 3.2 Data Analysis and Preparation
We first analyze each track’s training and testing data, as summarized in Tab. 1. Tracks 1 and 2 share the same training set, consisting of 23 sequences at 512$\times$512 resolution and 200 sequences at 640$\times$512 resolution. The training set for Track 3 is composed of 200 sequences at 640$\times$512 resolution. For testing, Tracks 1 and 2 each contain 216 non-overlapping sequences. Track 1’s test set is entirely at 640$\times$512 resolution, while Track 2 includes 16 sequences at 640$\times$512 and 200 at 512$\times$512. Track 3’s test set consists of 100 sequences at 640$\times$512 resolution. Additionally, Tab. 1 reports the width, height, and area distributions, along with their mean and standard deviation, providing essential insights for model hyperparameter tuning. Note that there may be slight differences in the numbers compared to the official release, as we have removed redundant annotations and defect cases, as illustrated in Fig. 1 (b).
@@
Table 1: Summary of data characteristics for training and testing. The training data for Tracks 1 and 2 are identical, while the testing data for Tracks 1 and 2 have no overlapping parts. Since Track 2 does not provide initial bounding boxes, we use "N/A" to indicate that this information is not applicable. Note that some bounding boxes were removed in Tracks 1 and 2 training sets due to labeled non-existent, zero-sized width or height, and cases where the boxes covered the entire image.
After analyzing the data, we split it for model training preparation. The number of frames and bounding boxes used for training, validation, and testing in the SOT and MOT tasks are detailed in Tab. 2. Specifically, Tracks 1 and 2 use YOLOv12 with BoT-SORT, while Track 3 employs YOLOv12 with BoT-SORT-ReID. Note that some numbers are in parentheses since we found the test set to provide limited information for the SOT task. Thus, the values in parentheses reflect the data only split into training and validation sets. Additionally, for BoT-SORT training, 1/10 of the data is primarily used to train the ReID module. This approach provides more effective ReID module training since many scenes are visually similar.
@@
Table 2: Data preparation summary for YOLOv12 and BoT-SORT-ReID across single-object (Tracks 1 and 2) and multi-object (Track 3) tracking. For YOLOv12 in SOT, numbers in parentheses indicate an alternative split with only training and validation sets. For BoT-SORT-ReID, numbers in parentheses indicate a reduced ReID training set selecting only the first 1/10 of frames from each sequence.
@@ 3.4 Training and Inference Strategies
The inference workflow is presented in Fig. 3. The overall procedure follows the original BoT-SORT scheme. However, we modify the output by reporting both online and lost targets for Tracks 1 and 2 while preserving the original output format for Track 3. We did not use linear track interpolation because ID switching frequently occurs due to camera motion or fast-moving UAVs, making interpolation ineffective for recovering missing detections. Instead, for the SOT task, we adopt a strategy based on the assumption that each frame contains at most one UAV, following this priority order: (1) report the UAV with the highest confidence score among the online targets, (2) if no online target is available, continue reporting the previous ID as the lost target in the subsequent tracker buffer frames, (3) if no previous ID is available, report the last known location until new online targets are detected. This strategy leverages the Kalman Filter’s prediction to accurately estimate the UAV’s location based on prior positions and velocity, significantly improving evaluation metrics in the SOT task. However, this strategy is not feasible for the MOT task due to the frequent overlap and ID switching between online and lost targets, which would lead to poor results. Therefore, we maintain the original output for Track 3 in this case.
Figure 3: YOLOv12n with BoT-SORT-SBS-S50 workflow diagram. The workflow follows the original BoT-SORT framework [1], with a slight revision: incorporating lost tracks to compensate for uninformative frames and improve object continuity. Specifically, for Tracks 1 and 2, lost target information is used to annotate potential object locations, while Track 3 retains the BoT-SORT original output.
@@ 4 Experimental Results
The experiments were conducted on two platforms: (1) a local system equipped with an Intel Core i7-12650H CPU, NVIDIA RTX 4050 GPU, and 16 GB of RAM, primarily used for data processing, and (2) a high-performance computing (HPC) system [27] featuring an NVIDIA H100 GPU and 80 GB of memory, used mainly for model training. All models were trained using the default settings (e.g., image input size of 640 and a track buffer of 30 frames) unless otherwise specified in the content or tables. This section begins by outlining the evaluation metrics for the three tracks, followed by the results for both the SOT and MOT tasks. We then present the leaderboard rankings and discuss key considerations and potential image enhancement techniques that could further improve UAV tracking.
@@ 4.1 Evaluation Metrics
Two evaluation metrics are used across the three competition tracks. The first metric applies to Tracks 1 and 2, where tracking accuracy is defined as:
@@
4.2 Evaluation Results on Tracks 1 and 2
@@
Table 3: Evaluation results for Tracks 1 and 2, summarizing eight trials per track. The first two rows detail YOLOv12 configurations, varying model sizes (n, s, m, l, x), and training epochs. The third row specifies the use of BoT-SORT ("✗" indicates exclusion; "✓" indicates inclusion). The fourth row lists the "min box area" threshold for filtering tiny bounding boxes, adjusted to 4 and 1 to account for tiny UAVs. The final row reports the resulting scores.
We present the evaluation results for Tracks 1 and 2 together, as both are SOT tasks, with the only difference being the presence of the initial UAV location. Eight meaningful trials are selected for both tracks, as shown in Tab. 3. Trials 1 and 2 serve as an ablation study to assess the impact of BoT-SORT. The results demonstrate a significant performance improvement: the score in Track 1 increases from 0.0786 to 0.5529, and in Track 2, it rises from 0.0992 to 0.3106 simply by adding BoT-SORT after the YOLOv12n detector. Trials 2 through 6 evaluate different detector model sizes (n, s, m, l, x), with the highest score achieved using YOLOv12l for both tracks. Trial 7 examines the effect of extended 300 epochs training, revealing a decline in performance compared to the 100 epochs training, likely due to overfitting. Finally, Trial 8 for each track shows the highest score we submitted, tuning the minimum box area threshold from 10 to 4 for Track 1 and from 10 to 1 for Track 2 to better capture smaller UAVs that may have been missed with the default setting.
@@ 4.5 Discussion and Enhancement Techniques
Figure 5: Potential frame enhancement techniques for multi-UAV tracking on a MultiUAV-262 video frame. From left to right: (1) original thermal infrared frame, (2) Sobel-based edge sharpening [10], (3) contrast enhancement via Contrast Limited Adaptive Histogram Equalization (CLAHE) [28], and (4) ReynoldsFlow+ visualization highlighting motion patterns to assist UAV detection [4].
@@
The evaluation results reveal several key insights. First, overfitting emerged due to our data-splitting strategy. To maximize scene diversity, we did not categorize videos by attributes such as fixed-camera setups or background types (e.g., sky or buildings). Instead, we directly split the dataset into training, validation, and testing sets, occasionally allowing frames from the same video to appear across splits. This likely contributed to overfitting, as evidenced by AP score discrepancies during local testing. Second, accurately rescaling the provided initial object positions to match the resolution used in training and inference is critical, as mismatches can mislead the tracker and degrade subsequent predictions. Third, increasing image resolution is key to breaking performance plateaus when parameter tuning fails to improve accuracy. For example, scaling from 640 to 1280 resolution yielded a significant score improvement of approximately 0.1. However, further increases produced diminishing gains as training at 2560 pixels for 7 epochs reached a score of 0.7072, and training at 3840 pixels for 1 epoch reached 0.7098, while both required significantly higher computational costs compared to training at 1280 pixels. Fourth, memory consumption during inference with YOLOv12 and BoT-SORT-ReID accumulates over time, leading to program crashes. To address this, we executed inference on a per-folder basis rather than processing all sequences in a single run. Finally, a clear performance gap is observed between runs with accurate initial object positions and those without, as evidenced by the performance difference between Tracks 1 and 2. This highlights the critical importance of promptly and reliably estimating initial positions to enhance tracking performance further.
@@
Additionally, as previously discussed, while our approach provides a strong baseline, it remains insufficient for achieving top-tier performance without further refinement. Fig. 4 (a) displays our model’s predictions across various scenarios, while Fig. 4 (b) highlights several key failure cases: (1) overlapping UAVs frequently cause ID switches, (2) distinguishing UAVs from flying creatures remains challenging, with the model often reassigning new IDs to UAVs following brief occlusions, (3) complex backgrounds lead to missed detections and tracking failures, and (4) tiny UAVs in cluttered environments provide little to no valuable information, making detection highly unreliable. The corresponding heatmaps in the last row illustrate the model’s inability to perceive UAVs effectively in these challenging conditions. These limitations emphasize the importance of image enhancement techniques to improve performance further. Fig. 5 illustrates several potential image enhancement methods. From left to right: (1) the original thermal infrared frame, (2) Sobel-based edge sharpening [10], which highlights edges more clearly than the original, (3) Contrast Limited Adaptive Histogram Equalization (CLAHE) [28], which improves contrast, and (4) ReynoldsFlow+ [4], a temporal enhancement method based on the Reynolds Transport Theorem [31], a three-dimensional generalization of the Leibniz integral rule [9], providing enhanced appearance for moving UAVs.